OpenStudy (moongazer):

Vector is 3m long, and vector is 4m long. The length of the sum of the vectors must be?

OpenStudy (anonymous):

vecotr is missing vector ID. vector unknown. vector line?

OpenStudy (amistre64):

between 4-3 and 4+3

OpenStudy (amistre64):



OpenStudy (moongazer):

how did you know that it is between 1 and 7?

OpenStudy (souvik):

what is the angle between those vectors?

OpenStudy (moongazer):

@souvik No angle was mentioned in the problem.

OpenStudy (souvik):

so the sum is maximum when the angle between them is zero degree and minimum when the angle is 180 degree.. the maximum sum =4+3=7 and minimum=4-3=1

OpenStudy (moongazer):

ok thanks! :)

OpenStudy (souvik):

you are welcome...:)

OpenStudy (moongazer):

Just a clarification, does it mean that you can't have an exact value for the sum because no angle or direction is specified?

OpenStudy (moongazer):

so you just determined what the value of the min and max (possible values) @souvik

OpenStudy (amistre64):

that is correct. Since the sum of vectors produces a visual "triangle"; then we can only assume that the third side is somewhere between 4-3 and 4+3 to account for all the possible lengths

OpenStudy (moongazer):

Thanks! :)